# The Multi-Query Paradox in Zeroth-Order Optimization

### v1

@@ 1 Introduction
Zeroth-order (ZO) optimization provides a powerful framework for problems where explicit gradient information is unavailable or prohibitively expensive to compute (Conn, Scheinberg, and Vicente 2009). Unlike first-order methods that rely on analytical gradients, ZO algorithms guide their search using only function value queries to approximate gradients. Typically, it is achieved by evaluating the objective function along one or more random directions and observing the change in function value (Spall 2002; Mania, Guy, and Recht 2018). This paradigm has recently gained significant attention due to its applicability to a wide array of challenging, high-dimensional problems, including fine-tuning of LLMs (Malladi et al. 2023; Zhang et al. 2024), policy optimization in reinforcement learning (Sun et al. 2020), and hyperparameter tuning (Li et al. 2021).
@@
The predominant approach in modern ZO optimization is the single-query method, where a gradient estimate is constructed using one random direction per optimization step (Nesterov and Spokoiny 2017). While foundational, these estimators are inherently noisy, with an estimation variance that often scales with the problem’s dimension (Nesterov and Spokoiny 2017). A natural extension to mitigate this issue is the multi-query approach, which employs a block of queries at each iteration to construct a more accurate gradient estimate (Duchi et al. 2015). However, this approach introduces a fundamental trade-off. Using multiple queries per iteration inverse proportionally reduces the total number of iteration steps that can be performed under a fixed budget of queries. A systematic investigation of the multi-query paradigm requires addressing two basic questions: (1) What is the most effective method for combining information from multiple query directions into a single, cohesive gradient estimate? (2) For a fixed total query budget, is it more efficient to perform many iterations with noisy, single-query estimates, or fewer iterations with higher-quality, multi-query estimates by consuming more queries per iteration?
@@
This work systematically addresses these questions. First, to resolve the aggregation problem, we propose a new multi-query estimator called Projection Alignment (ZO-Align), from the first principle of local surrogate minimization. While similar estimators have been proposed from the perspective of sparse gradient recovery (Cai et al. 2022; Wang et al. 2018), our optimization-centric derivation provides a more fundamental justification for the method, independent of any sparsity assumptions. For a comprehensive comparison, we also analyze the canonical strategy of simply averaging independent estimates (ZO-Avg) (Duchi et al. 2015). Our initial analysis of the estimators’ properties reveals that ZO-Align achieves a significantly lower mean squared error, motivating a deeper investigation into its performance within an optimization context.
@@
Second, to resolve the query allocation problem, we derive the convergence rates for algorithms using both ZO-Align and ZO-Avg, making the dependence on the number of queries used per iteration explicit. We then leverage these rates to formally pose and solve an optimization problem: for a fixed total query budget, find the best allocation of queries per iteration that minimizes the final optimization error. To ensure the generality and robustness of our conclusions, this entire theoretical framework is systematically applied across a wide spectrum of problem structures, including strongly convex, convex, non-convex, and stochastic settings.
Our analysis reveals a stark dichotomy where the optimal allocation strategy is critically dependent on the chosen aggregation method. For ZO-Avg, we prove that it cannot effectively utilize information from multiple queries; any attempt to use more than one query per iteration is detrimental, and the optimal strategy is always the single-query approach. In direct contrast, the ZO-Align estimator is capable of efficiently harnessing multi-query information. For ZO-Align, we prove that performance improves with the number of queries, making the optimal strategy a full-subspace estimation where the number of queries matches the problem’s dimension.
@@
To complement our theoretical analysis, we conducted extensive experiments across a suite of optimization problems. The empirical results consistently validate our theoretical findings, demonstrating the practical implications of the query allocation dichotomy. Our work clarifies that the choice is not about finding an optimal intermediate query size, but rather a commitment to an aggregation method that dictates whether the optimal algorithm is a sequential single-query method or a full-subspace estimation.
@@
  • We provide rigorous theoretical proofs that under a fixed query budget, the optimal query allocation for ZO-Avg is always single-query, whereas the optimal allocation for ZO-Align is to use the maximum possible number of queries, resolving the query allocation question for these two classes of estimators.
  • We empirically verify our theoretical findings, demonstrating that ZO-Align with a large query size significantly outperforms both ZO-Avg and single-query baselines across various optimization settings.
@@ 2 Related Work
Zeroth-order optimization approximates gradients using finite-difference techniques, a concept explored through various random perturbation strategies. For instance, theoretical convergence has been demonstrated using Gaussian random noise perturbations (Nesterov and Spokoiny 2017), while other approaches have utilized uniform sampling (Flaxman, Kalai, and McMahan 2004) and coordinate-wise perturbations (Lian et al. 2016). A distinct line of research focuses on problems where the gradient is assumed to be sparse, framing the estimation task as one of compressed sensing; here, techniques like LASSO (Wang et al. 2018) or other forms of regularized regression (Choromanski et al. 2020) enable the reconstruction of a sparse gradient from a limited number of function queries (Cai et al. 2022). More recently, methods based on linear interpolation have been investigated as a potentially more accurate alternative (Wang et al. 2024). Research suggests that linear interpolation can yield better gradient estimates than Gaussian smoothing (Berahas et al. 2019) and may require significantly fewer samples to do so (Berahas et al. 2022).
@@
A key challenge in ZOO is its high query complexity, motivating research into more efficient methods. Some methods build more accurate local models to estimate update direction by incorporating second-order Hessian information (Ye et al. 2025), or by leveraging problem geometry via Riemannian optimization (He et al. 2024). Other approaches optimize the querying process itself, such as using compressed sensing to achieve double-logarithmic query complexity for sparse problems (Qiu and Tong 2024), adaptively reusing old queries to reduce variance (Xiao, Ling, and Chen 2023), or learning a sampling policy with reinforcement learning (Zhai et al. 2024). However, despite these advances, these approaches are all centered on improving the quality of the estimate at each individual step. In contrast, the fundamental problem of how to optimally allocate a fixed query budget to maximize its utility remains a critical yet under-explored question.
@@ 3.1 Preliminaries and the Single-Query Estimator
We consider the unconstrained optimization problem $\min_{x\in\mathbb{R}^{d}}f(x)$ where direct access to the gradient $\nabla f(x)$ is unavailable.. Throughout this work, we assume the objective function $f$ is differentiable and $L$-smooth:
@@
The foundational building block for most ZO methods is the single-query estimator. It approximates the gradient by evaluating the function along a single random direction $u\in\mathbb{R}^{d}$. A common form is given by:
@@
where $\mu>0$ is a small smoothing parameter. This form is derived from the first-order Taylor expansion of $f(x)$ around $x$:
$f(x+\mu u)=f(x)+\mu u^{T}\nabla f(x)+O(\mu^{2}).$
For a sufficiently small $\mu$, the estimator approximates the projection of the true gradient onto the direction $u$: $\hat{g}_{u}(x)\approx(u^{T}\nabla f(x))u$. While fundamental, this single-direction estimate is inherently noisy, limiting its effectiveness. A natural extension is to probe the function along multiple directions to construct a more accurate gradient estimate.
@@ 3.2 Simple Averaging Estimator (ZO-Avg)
$\hat{g}_{\text{AVG}}(x)=\frac{1}{q}\sum_{i=1}^{q}\frac{f(x+\mu u_{i})-f(x)}{\mu}u_{i}.$
Our subsequent theoretical analysis focuses on the idealized estimators where the bias from the finite difference is negligible (i.e., as $\mu\to 0$). This allows us to isolate the error stemming from the geometric properties of the estimators. In this idealized form, the estimator approximates:
$\hat{g}_{\text{AVG}}(x)\approx\left(\frac{1}{q}\sum_{i=1}^{q}u_{i}u_{i}^{T}\right)\nabla f(x)=\frac{1}{q}UU^{T}\nabla f(x).$
This estimator is appealing due to its simplicity and straightforward parallel implementation. A key property of ZO-Avg is that it provides an unbiased estimate of the true gradient.
Proposition 3.1.
Let the columns of $U$ be i.i.d. samples from $\mathcal{N}(0,I)$. The ZO-Avg estimator is unbiased, i.e., $\mathbb{E}[\hat{g}_{AVG}(x)]=\nabla f(x)$.
Proof.
See Appendix A.1. ∎
While unbiasedness is a desirable property, it does not guarantee efficiency. As we will show, the high variance of this estimator plays a critical role in its convergence behavior.
@@ 3.3 Projection Alignment Method (ZO-Align)
$\hat{g}_{{\text{ALG}}}(x)=U(U^{T}U)^{-1}\begin{bmatrix}\frac{f(x+\mu u_{1})-f(x)}{\mu}\\ \vdots\\ \frac{f(x+\mu u_{q})-f(x)}{\mu}\end{bmatrix}$.
@@
Unlike ZO-Avg, ZO-Align is a biased estimator.
Proposition 3.2.
Let the columns of $U$ be i.i.d. samples from $\mathcal{N}(0,I)$. The ZO-Align estimator has an expected value of $\mathbb{E}[\hat{g}_{ALG}(x)]=\frac{q}{d}\nabla f(x)$.
Proof.
See Appendix A.1. ∎
@@ 3.4 Accuracy Analysis: A Tale of Two Estimators
The two estimators present a classic statistical trade-off. ZO-Avg is unbiased but suffers from high variance. In contrast, ZO-Align accepts a well-structured bias in exchange for a dramatic reduction in estimation error. To formalize this comparison, we analyze the Mean Squared Error (MSE) of each estimator, defined as $\mathbb{E}[\|\hat{g}(x)-\nabla f(x)\|^{2}]$, which captures the total error from both bias and variance.
@@
See Appendix A.1. ∎
For any $q>1$, it is clear that $\frac{d-q}{d}\ll\frac{d+1}{q}$, indicating that ZO-Align offers a significantly lower estimation error. This fundamental difference in estimator quality is the primary driver of the performance gap we will establish in the following sections. It highlights that the principled, model-based approach of ZO-Align is far more effective at aggregating information from multiple queries than the naive averaging of ZO-Avg.
@@ 4.1 Query Allocation Problem
Suppose a fixed total query budget, $K$, an algorithm performs a series of updates over $T$ iterations, and at each iteration $t$, it uses a set of $q_{t}$ function queries as stated in Section 3.2. The total number of queries is therefore constrained by $\sum_{t=0}^{T-1}q_{t}\leq K$. Our central goal is to determine the optimal query allocation strategy—the sequence of queries per iteration, $\{q_{t}\}_{t=0}^{T-1}$—that minimizes the final optimization error.
@@ 4.2 Optimal Query Allocation
We begin by analyzing the case where function evaluations are deterministic. The update rule at iteration $t$ is $x_{t+1}=x_{t}-\eta_{t}\hat{g}(x_{t})$, where $\hat{g}$ is either the ZO-Avg or ZO-Align estimator constructed with a set of $q_{t}$ query directions, as illustrated in Section 3.
Strongly Convex Case
A function is defined as $\mu$-strongly convex if there exists a constant $\mu$ such that for all $x$ and $y$ in its domain:
$f(y)\geq f(x)+\nabla f(x)^{T}(y-x)+\frac{\mu}{2}\|y-x\|^{2}.$
When the objective function $f(x)$ is both $L$-smooth and $\mu$-strongly convex, we can achieve a linear convergence rate (Boyd and Vandenberghe 2004). However, the rate’s dependence on the number of queries per iteration, $q_{t}$, reveals the core dichotomy of our work. We first analyze the ZO-Avg estimator.
@@
Let $f$ be an $L$-smooth and $\mu$-strongly convex function. The ZO-Avg algorithm with the optimal step size $\eta_{t}=\frac{q_{t}}{L(q_{t}+d+1)}$ achieves the following convergence guarantee:
$\begin{split}&\quad\mathbb{E}[f(x_{T})-f(x^{*})]\\ \leq&\left(\prod_{t=0}^{T-1}\left(1-\frac{\mu q_{t}}{L(q_{t}+d+1)}\right)\right)(f(x_{0})-f(x^{*})),\end{split}$
@@
$\mathbb{E}[f(x_{K})-f(x^{*})]\leq\left(1-\frac{\mu}{L(d+2)}\right)^{K}(f(x_{0})-f(x^{*}))$
@@
See Appendix A.1. ∎
The result in Theorem 4.1 is deeply counter-intuitive. It suggests that when using the simple averaging estimator, allocating more than one query per iteration is detrimental to the convergence rate. The core reason lies in the trade-off between variance reduction and the cost of queries. While increasing $q_{t}$ does reduce the variance of the gradient estimate, the improvement scales as $1/q_{t}$ as shown in Section 3.4. However, the cost increases linearly with $q_{t}$. The analysis shows that the marginal benefit of reducing variance is not worth the cost of the additional queries. It is always better to spend the query budget on taking more, albeit noisier, steps. This exposes a fundamental flaw in the simple average strategy: it is an inefficient use of a limited query budget.
@@
Let $f$ be an $L$-smooth and $\mu$-strongly convex function. The ZO-Align algorithm with the optimal step size $\eta_{t}=1/L$ achieves the following convergence guarantee:
$\mathbb{E}[f(x_{T})-f(x^{*})]\leq\left(\prod_{t=0}^{T-1}\left(1-\frac{\mu q_{t}}{Ld}\right)\right)(f(x_{0})-f(x^{*})),$
@@
$\mathbb{E}[f(x_{K/d})-f(x^{*})]\leq\left(1-\frac{\mu}{L}\right)^{K/d}(f(x_{0})-f(x^{*})).$
@@
See Appendix A.1. ∎
Theorem 4.2 presents a picture opposite to that of ZO-Avg. For ZO-Align, the optimal strategy is to use the maximum number of queries per step. The intuition is that ZO-Align does not simply average away noise; it constructs a principled, low-error projection of the gradient onto the query subspace. As shown in Section 3.4, the quality of this projection improves substantially as the number of query directions increases. Each additional query provides a significant marginal benefit by better aligning the search direction with the true gradient. It is therefore more effective to spend the budget forming a single high-quality, confident step direction than to take many small, uncertain steps. This result highlights the power of a model-based approach: by aggregating information correctly, investing more queries per step becomes a powerful tool for accelerating convergence.
Convex Case

### v3

@@ 1 Introduction
Zeroth-order (ZO) optimization provides a powerful framework for problems where explicit gradient information is unavailable or prohibitively expensive to compute (conn2009introduction). Unlike first-order methods that rely on analytical gradients, ZO algorithms guide their search using only function value queries to approximate gradients. Typically, it is achieved by evaluating the objective function along one or more random directions and observing the change in function value (spall2002multivariate; mania2018simple). This paradigm has recently gained significant attention due to its applicability to a wide array of challenging, high-dimensional problems, including fine-tuning of LLMs (malladi2023fine; zhang2024revisiting), policy optimization in reinforcement learning (sun2020zeroth), and hyperparameter tuning (li2021zeroth).
@@
The predominant approach in modern ZO optimization is the single-query method, where a gradient estimate is constructed using one random direction per optimization step (nesterov2017random). While foundational, these estimators are inherently noisy, with an estimation variance that often scales with the problem’s dimension (nesterov2017random). A natural extension to mitigate this issue is the multi-query approach, which employs a block of queries at each iteration to construct a more accurate gradient estimate (duchi2015optimal).
However, this extension introduces a fundamental trade-off. Using multiple queries per iteration proportionally reduces the total number of iterations that can be performed under a fixed query budget. This tension gives rise to what we call the multi-query paradox: increasing the number of queries per iteration improves the quality of each gradient estimate but simultaneously decreases the total number of updates one can afford. A systematic investigation of this paradox requires addressing two basic questions: (1) What is the most effective method for combining information from multiple query directions into a single, coherent gradient estimate? (2) For a fixed total query budget, is it more efficient to perform many iterations with noisy, single-query estimates, or fewer iterations with higher-quality, multi-query estimates that consume more queries per iteration?
@@
This work systematically addresses these questions. First, to resolve the aggregation problem, we propose a new multi-query estimator called Projection Alignment (ZO-Align), derived from first principles via local surrogate minimization. While similar estimators have been proposed from the perspective of sparse gradient recovery (cai2022zeroth; wang2018stochastic), our optimization-centric derivation provides a more fundamental justification for the method, independent of any sparsity assumptions. For a comprehensive comparison, we also analyze the canonical strategy of simply averaging estimates (ZO-Avg) (duchi2015optimal). Our initial analysis of the estimators’ properties reveals that ZO-Align achieves a significantly lower mean squared error, motivating a deeper investigation into its performance within an optimization context.
@@
Second, to resolve the query allocation problem, we derive convergence rates for algorithms using both ZO-Align and ZO-Avg, making the dependence on the number of queries used per iteration explicit. We then leverage these rates to formally pose and solve an optimization problem: given a fixed total query budget, determine the allocation of queries per iteration that minimizes the final optimization error. To ensure the generality and robustness of our conclusions, this theoretical framework is systematically applied across a broad spectrum of problem structures, including strongly convex, convex, non-convex, and stochastic settings.
Our analysis reveals a stark dichotomy in optimal allocation strategies, critically dependent on the chosen aggregation method. For ZO-Avg, we prove that it cannot effectively utilize information from multiple queries: under a fixed budget, any attempt to use more than one query per iteration is detrimental, and the optimal strategy is always the single-query approach. In contrast, the ZO-Align estimator combines multi-query information in a geometry-aware way that avoids wasting queries on redundant directions, and in strongly convex and stochastic problems multi-query ZO-Align is more query-efficient than any single-query scheme. We further analyze the regime where the ambient dimension is much larger than the query block size, showing that ZO-Avg continues to follow the same single-query allocation rule, while ZO-Align degenerates into a step direction that is nearly colinear with a suitably rescaled ZO-Avg update due to the near-orthogonality of random directions in high dimensions.
@@
To complement our theoretical analysis, we conducted extensive experiments across a suite of optimization problems. The empirical results consistently validate our theoretical findings, illustrating the practical implications of the query allocation dichotomy. Our work clarifies that the central design choice is not about tuning an intermediate query size, but rather about committing to an aggregation method that dictates whether the optimal algorithm is a sequential single-query method or a full-subspace estimation procedure.
@@
  • We provide rigorous theoretical proofs that under a fixed query budget, the optimal query allocation for ZO-Avg is always single-query, whereas ZO-Align can make better use of multi-query information, resolving the query allocation question for these two classes of estimators.
  • We empirically validate our theoretical findings across a range of optimization problems, and also include high-dimensional tasks that illustrates how our query-allocation insights manifest in large-scale problems.
@@ 2 Related Work
Zeroth-order optimization approximates gradients using finite-difference techniques, a concept explored through various random perturbation strategies. For instance, theoretical convergence has been demonstrated using Gaussian random noise perturbations (nesterov2017random), while other approaches have utilized uniform sampling (flaxman2004online) and coordinate-wise perturbations (lian2016comprehensive). A distinct line of research focuses on problems where the gradient is assumed to be sparse, framing the estimation task as one of compressed sensing; here, techniques like LASSO (wang2018stochastic) or other forms of regularized regression (choromanski2020provably) enable the reconstruction of a sparse gradient from a limited number of function queries (cai2022zeroth).
@@
A central challenge in zeroth-order optimization is its high query complexity, which has motivated a broad range of techniques aimed at improving query efficiency. Some methods construct more accurate local models for the update direction by incorporating second-order information (ye2025hessian) or by exploiting problem geometry via Riemannian optimization (he2024riemannian). Other approaches focus on optimizing the querying process itself, for example by using compressed sensing to achieve double-logarithmic query complexity in sparse settings (qiu2024gradient), adaptively reusing past queries to reduce variance (xiao2023lazy), or learning query directions through reinforcement learning (zhai2024learning). Despite their differences, these approaches share a common focus: they seek to improve the quality of the gradient estimate at each individual iteration, typically under a fixed number of queries per step.
In contrast, our work takes a complementary perspective. Rather than designing yet another estimator for a given per-iteration query budget, we ask how to optimally allocate a fixed total query budget across iterations and aggregation schemes. Our analysis shows that, even for two canonical multi-query estimators, this allocation problem exhibits a sharp dichotomy: the optimal query-per-iteration choice collapses to opposite extremes depending on the aggregation method used.
@@ 3.1 Preliminaries and the Single-Query Estimator
We consider the unconstrained optimization problem $\min_{x\in\mathbb{R}^{d}}f(x)$ where direct access to the gradient $\nabla f(x)$ is unavailable. Throughout this work, we assume the objective function $f$ is differentiable and $L$-smooth:
@@
The foundational building block for most ZO methods is the single-query estimator. It approximates the gradient by evaluating the function along a single random direction $u\in\mathbb{R}^{d}$ using a one-sided finite-difference scheme. A common form is given by:
@@
where $\mu>0$ is a small smoothing parameter. Throughout this work, we focus on an idealized regime in which $\mu\to 0$ and the estimator approximates the projection of the true gradient onto the direction $u$:
$\hat{g}_{u}(x)\overset{\mu\to 0}{\approx}(u^{\top}\nabla f(x))u.$
All MSE and convergence results below are derived under this idealization. Importantly, the same geometric analysis applies to two-sided finite-difference estimators, which differ from the one-sided form only in the higher-order $\mu$-dependent terms (nesterov2017random).
While fundamental, this single-direction estimate is inherently noisy, limiting its effectiveness. A natural extension is to probe the function along multiple directions to construct a more accurate gradient estimate.
@@ 3.2 Simple Averaging Estimator (ZO-Avg)
$\hat{g}_{\text{AVG}}(x)$ $=\frac{1}{q}\sum_{i=1}^{q}\frac{f(x+\mu u_{i})-f(x)}{\mu}u_{i}$ (1)
$\overset{\mu\to 0}{\approx}\left(\frac{1}{q}\sum_{i=1}^{q}u_{i}u_{i}^{T}\right)\nabla f(x)=\frac{1}{q}UU^{T}\nabla f(x).$
This estimator is appealing due to its simplicity and straightforward parallel implementation.
@@ 3.3 Projection Alignment Method (ZO-Align)
$\hat{g}_{{\text{ALG}}}(x)=U(U^{T}U)^{-1}\begin{bmatrix}\frac{f(x+\mu u_{1})-f(x)}{\mu}\\ \vdots\\ \frac{f(x+\mu u_{q})-f(x)}{\mu}\end{bmatrix}$. (2)
@@ 3.4 Accuracy Analysis: A Tale of Two Estimators
To formalize this comparison, we analyze the MSE of each estimator, defined as $\mathbb{E}[\|\hat{g}(x)-\nabla f(x)\|^{2}]$, which captures the total error from both bias and variance.
@@
See Appendix B.3. ∎
For any $q>1$, it is clear that $\frac{d-q}{d}<\frac{d+1}{q}$, indicating that ZO-Align offers a significantly lower estimation error. This fundamental difference in estimator quality is the primary driver of the performance gap we will establish in the following sections. It highlights that the principled, model-based approach of ZO-Align is far more effective at aggregating information from multiple queries than the naive averaging of ZO-Avg.
@@ 4.1 Query Allocation Problem
Suppose we are given a fixed total query budget $K$. An algorithm performs a series of updates over $T$ iterations, and at each iteration $t$ it uses a set of $q_{t}$ function queries as stated in Section 3.2. The total number of queries is therefore constrained by $\sum_{t=0}^{T-1}q_{t}\leq K$. Our central goal is to determine the optimal query allocation strategy—the sequence of queries per iteration, $\{q_{t}\}_{t=0}^{T-1}$—that minimizes the final optimization error.
@@ 4.2 Optimal Query Allocation
We begin by analyzing the case where function evaluations are deterministic. The update rule at iteration $t$ is $x_{t+1}=x_{t}-\eta_{t}\hat{g}(x_{t})$, where $\hat{g}$ is either the ZO-Avg or ZO-Align estimator constructed with a set of $q_{t}$ query directions, as illustrated in Section 3. Given an estimator with $q_{t}$ queries at iteration $t$, we choose the step size $\eta_{t}$ by minimizing the standard $L$-smooth upper bound on the expected decrease of $f$. The following proposition summarizes the resulting choices; these step sizes will be used throughout this section.
Proposition 4.1.
Let $f$ be an $L$-smooth function. Consider a single update of the form $x^{+}=x-\eta\hat{g}(x)$. In the idealized regime $\mu\to 0$, the following holds:
(ZO-Avg) If $\hat{g}=\hat{g}_{\text{AVG}}$ is the averaging estimator in (1), then the step size that maximizes the guaranteed decrease is
$\eta_{\text{AVG}}^{*}(q)=\frac{q}{L(q+d+1)}.$ (3)
(ZO-Align) If $\hat{g}=\hat{g}_{\text{ALG}}$ is the projection-alignment estimator in (2), then step size that maximizes the guaranteed decrease in this bound is
$\eta_{\text{ALG}}^{*}(q)=\frac{1}{L}.$ (4)
Proof.
See Appendix B.4. ∎
4.2.1 Strongly Convex Case
A function is defined as $\gamma$-strongly convex if there exists a constant $\gamma$ such that for all $x$ and $y$ in its domain:
$f(y)\geq f(x)+\nabla f(x)^{T}(y-x)+\frac{\gamma}{2}\|y-x\|^{2}.$
When the objective function $f(x)$ is both $L$-smooth and $\gamma$-strongly convex, we can achieve a linear convergence rate (boyd2004convex). However, the rate’s dependence on the number of queries per iteration, $q_{t}$, reveals the core dichotomy of our work. We first analyze the ZO-Avg estimator.
@@ 4.2.1 Strongly Convex Case
Let $f$ be an $L$-smooth and $\gamma$-strongly convex function. The ZO-Avg algorithm with the optimal step size stated in (3) achieves the following convergence guarantee:
$\begin{split}&\quad\mathbb{E}[f(x_{T})-f(x^{*})]\\ \leq&\left(\prod_{t=0}^{T-1}\left(1-\frac{\gamma q_{t}}{L(q_{t}+d+1)}\right)\right)(f(x_{0})-f(x^{*})),\end{split}$
@@
$\mathbb{E}[f(x_{K})-f(x^{*})]\leq\left(1-\frac{\gamma}{L(d+2)}\right)^{K}(f(x_{0})-f(x^{*}))$
@@
See Appendix B.5. ∎
The result in Theorem 4.2 suggests that when using the simple averaging estimator, allocating more than one query per iteration is detrimental to the convergence rate. The core reason lies in the trade-off between variance reduction and the cost of queries. While increasing $q_{t}$ does reduce the variance of the gradient estimate, the improvement scales sublinearly $1/q_{t}$, as shown in Section 3.4. However, the cost increases linearly with $q_{t}$. The analysis shows that the marginal benefit of reducing variance is not worth the cost of the additional queries. It is always better to spend the query budget on taking more, albeit noisier, steps. This exposes a fundamental flaw in the simple average strategy: it is an inefficient use of a limited query budget.
@@
Let $f$ be an $L$-smooth and $\gamma$-strongly convex function. The ZO-Align algorithm with the optimal step size $\eta_{t}=1/L$ achieves the following convergence guarantee:
$\mathbb{E}[f(x_{T})-f(x^{*})]\leq\left(\prod_{t=0}^{T-1}\left(1-\frac{\gamma q_{t}}{Ld}\right)\right)(f(x_{0})-f(x^{*})),$
@@
$\mathbb{E}[f(x_{K/d})-f(x^{*})]\leq\left(1-\frac{\gamma}{L}\right)^{K/d}(f(x_{0})-f(x^{*})).$
@@
See Appendix B.6. ∎
For ZO-Align, the optimal strategy is to use the maximum number of queries per step. The intuition is that ZO-Align does not simply average away noise; it constructs a principled, low-error projection of the gradient onto the query subspace. Optimization error is reduced multiplicatively at each step, meaning a more accurate gradient provides an exponential advantage over time. It is therefore more effective to spend the budget forming a single high-quality step than to take many uncertain steps.
4.2.2 Convex Case